If a 12 percent rise in the price of orange juice decreases the quantity of orange juice demanded


by 22 percent and increases the quantity of apple juice demanded by 14 percent, calculate the


a. Price elasticity of demand for orange juice.


b. Cross elasticity of demand for apple juice with respect to the price of orange juice.

1
Expert's answer
2022-01-05T11:14:15-0500

Solution:

a.). Price elasticity of demand for orange juice:

Price elasticity of demand for orange juice = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"

= "\frac{-22\% }{12\% } = -1.8"

Price elasticity of demand for orange juice = 1.8

 

b.). Cross elasticity of demand for apple juice to the price of orange juice:

Cross elasticity of demand for apple juice to the price of orange juice = "=\frac{\%\;change\; in\; quantity\; demanded\; for\; apple\; juice}{\%\; change\; in\; price\; of\; orange\; juice}"


= "\frac{14\% }{12\% } = 1.2"

 

Cross elasticity of demand for apple juice to the price of orange juice = 1.2